### v1

# Learning Getting-Up Policies for Real-World Humanoid Robots

Abstract
Automatic fall recovery is a crucial prerequisite before humanoid robots can be reliably deployed. Hand-designing controllers for getting up is difficult because of the varied configurations a humanoid can end up in after a fall and the challenging terrains humanoid robots are expected to operate on. This paper develops a learning framework to produce controllers that enable humanoid robots to get up from varying configurations on varying terrains. Unlike previous successful applications of humanoid locomotion learning, the getting-up task involves complex contact patterns, which necessitates accurately modeling the collision geometry and sparser rewards. We address these challenges through a two-phase approach that follows a curriculum. The first stage focuses on discovering a good getting-up trajectory under minimal constraints on smoothness or speed / torque limits. The second stage then refines the discovered motions into deployable (i.e. smooth and slow) motions that are robust to variations in initial configuration and terrains. We find these innovations enable a real-world G1 humanoid robot to get up from two main situations that we considered: a) lying face up and b) lying face down, both tested on flat, deformable, slippery surfaces and slopes (e.g., sloppy grass and snowfield). To the best of our knowledge, this is the first successful demonstration of learned getting-up policies for human-sized humanoid robots in the real world. Project page: https://humanoid-getup.github.io/
Figure 1: HumanUP provides a simple and general two-stage training method for humanoid getting-up tasks, which can be directly deployed on Unitree G1 humanoid robots [70]. Our policies showcase robust and smooth behavior that can get up from diverse lying postures (both supine and prone) on varied terrains such as grass slopes and stone tile.
00footnotetext: Equal contributions.
I Introduction
This paper develops learned controllers that enable a humanoid robot to get up from varied fall configurations on varied terrains. Humanoid robots are susceptible to falls, and their reliance on humans for fall recovery hinders their deployment. Furthermore, as humanoid robots are expected to work in environments involving complex terrains and tight workspaces (i.e. challenging scenarios that are too difficult for wheeled robots), a humanoid robot may end up in an unpredictable configuration upon a fall, or may be on an unknown terrain. 26 of the 46 trials at the DARPA Robotics Challenge (DRC) had a fall and 25 of these falls required human intervention for recovery [40]. The DRC identified fall prevention and recovery as a major topic needing more research. This paper pursues this research and proposes a learning-based framework for learning fall recovery policies for humanoid robots under varying conditions.
The need for recovering from varied initial conditions makes it hard to design a fall recovery controller by hand and motivates the need for learning via trial and error in simulation. Such learning has produced exciting results in recent years for locomotion problems involving quadrupeds and humanoids, e.g. [43, 60]. Motivated by these exciting results, we started with simply applying the Sim-to-Real (Sim2Real) paradigm for the getting-up problem. However, we quickly realized that the getting-up problem is different from typical locomotion problems in the following three significant ways that made a naive adaptation of previous work inadequate:
  • a)
    Non-periodic behavior. In locomotion, contacts with the environment happen in structured ways: cyclic left-right stepping pattern. The getting-up problem doesn’t have such a periodic behavior. The contact sequence necessary for getting up itself needs to be figured out. This makes optimization harder and may render phase coupling of left and right feet commonly used in locomotion ineffective.
  • b)
    Richness in contact. Different from locomotion, contacts necessary for getting up are not limited to just the feet. Many other parts of the robot are likely already in touch with the terrain. But more importantly, the robot may find it useful to employ its body, outside of the feet, to exert forces upon the environment, in order to get up. Freezing / decoupling the upper body, only coarsely modeling the upper body for collisions, and using a larger simulation step size: the typical design choices made in locomotion, are no longer applicable for the getting up task.
  • c)
    Reward sparsity. Designing rewards for getting up is harder than other locomotion tasks. Velocity tracking offers a dense reward and feedback on whether the robot is meaningfully walking forward is available within a few tens of simulation steps. In contrast, many parts of the body make negative progress, e.g., the torso first needs to tilt down for seconds before tilting up to finally get up.
We present HumanUP, which circumvents these problems via a two-stage reinforcement learning (RL) based training. Stage I targets task solving in easier settings (sparse task rewards with weak regularization), while Stage II solves the task of making the learned motion deployable (i.e., control should be smooth; velocities and executed torques should be small; etc). To be specific, Stage I is designed for motion discovery without being limited by smoothness in motion or speed / torque limits, which tackles a hard task under sparse and under-specified task rewards while being less constrained for ensuring task solving. Stage II is optimized to track the state trajectory discovered in the first stage to tackle easier motion tracking with dense tracking rewards, which is under strict Sim2Real control regularization for ensuring Sim2Real transfer. From Stage I to Stage II, we employ a Sim2Real learning curriculum that progresses from simplified $\rightarrow$ full collision mesh, canonical $\rightarrow$ random initial lying posture, and weak to strong control regularization and domain randomization. This two-stage approach integrates a hard-to-easy task-solving curriculum with an easy-to-hard Sim2Real curriculum, both of which are crucial for successful learning, as demonstrated in our experiments.
We conduct experiments in simulation and the real world with the G1 platform from Unitree. In the real world, we find our framework enables the G1 robot to get up from three situations that we considered: a) supine (lying face up) poses and b) prone (lying face down) poses. This expands the capability that the G1 comes with: the default hand-crafted getting-up controller only successfully gets up from supine poses on a flat surface without bumps. In simulated experiments, we find that our framework can successfully learn getting-up policies that work on varied terrains, varied starting poses.
II Related Work
We review related works on humanoid control, learning for humanoid control, and work specifically targeted toward fall recovery for legged robots.
II-A Humanoid Control
Controlling a high degree of freedom humanoid robots have fascinated researchers for the last several decades. Model-based techniques, such as those based on the Zero Moment Point (ZMP) principle [71, 31, 61, 69], optimization [41, 4, 14], and Model Predictive Control (MPC) [12, 21, 15, 74], have demonstrated remarkable success in fundamental locomotion tasks like walking, running and jumping. However, these approaches often struggle to generalize or adapt to novel environments. In contrast, learning-based approaches have recently made significant strides, continuously expanding the generalization capabilities of humanoid locomotion controllers.
Figure 2: HumanUP system overview. Our getting-up policy (Section III-A) is trained in simulation using two-stage RL training, after which it is directly deployed in the real world. (a) Stage I (Section III-B1) learns a discovery policy $f$ that figures out a getting-up trajectory with minimal deployment constraints. (b) Stage II (Section III-B2) converts the trajectory discovered by Stage I into a policy $\pi$ that is deployable, robust, and generalizable. This policy $\pi$ is trained by learning to track a slowed down version of the discovered trajectory under strong control regularization on varied terrains and from varied initial poses. (c) The two-stage training induces a curriculum (Section III-C). Stage I targets motion discovery in easier settings (simpler collision geometry, same starting poses, weak regularization, no variations in terrain), while Stage II solves the task of making the learned motion deployable and generalizable.
II-A1 Learning for humanoid control
Learning in simulation via reinforcement followed by a sim-to-real transfer has led to many successful locomotion results for quadrupeds [43, 42] and humanoids [60, 59, 58, 26, 8, 2]. This has enabled locomotion on challenging in-the-wild terrain [58, 25], agile motions like jumping [44, 76], and even locomotion driven by visual inputs [77, 46]. Researchers have also expanded the repertoire of humanoid motions to skillful movements like dancing and naturalistic walking gaits through use of human mocap or video data [34, 9, 53, 30]. Some works address locomotion and manipulation problems for humanoids simultaneously to enable loco-manipulation controllers in an end-to-end fashion facilitated by teleportation [29, 19, 48]. Notably, these tasks mostly involve contact between the feet and the environment, thus requiring only limited contact reasoning. How to effectively develop controllers for more contact-rich tasks like crawling, tumbling, and getting up that require numerous, dynamic, and unpredictable contacts between the whole body and the environment remains under-explored.
II-B Legged robots fall recovery
Humanoid robots are vulnerable to falls due to under-actuated control dynamics, high-dimensional states, and unstructured environments [38, 31, 24, 40, 32, 27], making the ability to recover from falling of great significance. Over the years, this problem has been tackled in the following ways.
II-B1 Getting up via motion planning
Early work from Morimoto and Doya [54] solved the getting-up problem for a two-joint, three-link walking robot in 2D, and several discrete states are used as subgoals to transit via hierarchical RL. This line of work can be viewed as an application of motion planning by configuration graph transition learning [39], where stored robot states between lying and standing are used as graph nodes to transit [20, 36, 37, 64]. More recently, some progress has been made to enable toy-sized humanoid robots to get up. For example, González-Fierro et al. [23] explores getting up from a canonical sitting posture with motion planning by imitating human demonstration with ZMP ceriterion. To address the high-dimensionality of humanoid configurations, Jeong and Lee [33] leverage bilateral symmetry to reduce the control DoFs by half and a clustering technique is used for further reducing the complexity of configuration space, thereby improving getting-up learning efficiency. However, such state machine learning using predefined configuration graphs may not be sufficient for generalizing to unpredictable initial and intermediate states, which happens when the robot operates on challenging terrains.
II-B2 Hand-designed getting-up trajectories
Another solution, often adopted by commercial products, is to replay a manually designed motion trajectory. For example, Unitree [70] has a getting-up controller built into G1’s default controllers. Booster Robotics [1] designed a specific recovery controller for their robots that can help the robot recover from fallen states. The main drawback of such pre-defined trajectory getting-up controllers is that they can only handle a limited number of fallen states and lack generalization, as our experimental comparisons will show.
II-B3 Learned getting-up policies for real robots
RL followed by sim-to-real has also been successfully applied for quadruped [43, 35, 72, 51] fall recovery. For example, Lee et al. [43] explore sim2real RL to achieve real-world quadruped fall recovery from complex configurations. Ji et al. [35] train a recovery policy that enables the quadruped to dribble in snowy and rough terrains continuously. Wang et al. [72] develop a quadruped recovery policy in highly dynamic scenarios..
II-B4 Learned getting-up policies for character animation
A parallel research effort in character animation, also explores the design of RL-based motion imitation algorithms: DeepMimic [55], AMP [56], PHC [49], among others [11, 50, 67, 5, 22, 73]. These have also demonstrated successful getting-up controllers in simulation. By tracking user-specified getting-up curves, Frezzato et al. [17] enable humanoid characters to get up by synthesizing physically plausible motion. Without recourse to mocap data, such naturalistic getting-up controllers for simulated humanoid characters can also be developed with careful curriculum designs [65]. Some works explore sampling-based methods for addressing contact-rich character locomotion including getting up [45, 28, 57], while some works have demonstrated success in humanoid getting up with online model-predictive control [66]. It is worth noticing, however, that these works use humanoid characters with larger DoFs compared to humanoid robots (e.g., 69 DoFs in SMPL [47]) and use simplified dynamics. As a result, learned policies operate body parts at high velocities and in infeasible ways, leading to behavior that cannot be transferred into the real world directly. Hence, developing generalizable recovery controllers for humanoid robots remains an open problem.
III HumanUP: Sim-to-Real Humanoid Getting Up
Our goal is to learn a getting-up policy $\pi$ that enables a humanoid to get up from arbitrary initial postures. We consider getting up from two families of lying postures: a) supine poses (i.e. lying facing up) and b) prone poses (i.e. lying face down). Getting up from these two groups of postures may require different behaviors, which makes it challenging to learn a single policy that handles both scenarios. To tackle this issue, we decompose the getting-up task from a prone pose to first rolling over and then standing up from the resulting supine posture. Therefore, our objective is to learn policies for rolling over from a prone pose and getting up from a supine pose separately.
To solve these two tasks, we propose HumanUP, a general learning framework for training getting-up and rolling over policies, which is illustrated in Fig. 2. In stage I, a discovery policy $f$ is trained to figure out standing-up or rolling-over motions. $f$ is trained without deployment constraints, and only our task and symmetry rewards are used. In stage II, a deployable policy $\pi$ imitates the rolling-over / getting-up behaviors obtained from stage I under strong control regularization. This deployable policy $\pi$ is transferred from simulation to the real world as the final policy. To overcome the difficulty of contact-rich locomotion learning involved in these tasks, we design a curricula for learning. We detail the policy model and two-stage training in Section III-B, and the curriculum design is introduced in Section III-C.
III-A Policy Architecture
HumanUP trains two policy models $f$ and $\pi$ with RL. Both policy models take observation $\mathbf{o}_{t}=[\mathbf{z}_{t},\mathbf{s}_{t},\mathbf{s}_{t-10:t-1}]\in\mathbb{R}^{868}$ as input and output action $\mathbf{a}_{t}\in\mathbb{R}^{23}$, where $\mathbf{s}_{t}\in\mathbb{R}^{74}$ is the robot’s proprioceptive information, $\mathbf{s}_{t-10:t-1}$ is the $10$ steps history states, and $\mathbf{z}_{t}\in\mathbb{R}^{54}$ are the encoded environment extrinsic latents learned using regularized online adaptation [18]. We use the proprioceptive information $\mathbf{s}_{t}$ consisting of the robot’s roll and pitch, angular velocity, DoF velocities, and DoF positions. Such proprioceptive information can be accurately obtained in the real world, and we find that these are sufficient for the robot to infer the overall posture. We do not use any linear velocity and yaw information as it is difficult to reliably estimate them in the real world [29].
The policy models are implemented as MLPs and trained via PPO [62]. The optimization targets maximizing the expected $\gamma$-discounted policy return within $T$ episode length: $\mathbb{E}\left[\sum_{t=1}^{T}\gamma^{t-1}r_{t}\right]$, where $r_{t}$ is the reward at timestamp $t$.
III-B Two-Stage Policy Learning
III-B1 Stage I: Discovery Policy
This stage discovers getting-up / rolling-over behavior efficiently without deployment constraints. We use the following task rewards with very weak regularization to train this discovery policy $f$. We describe the rewards used for optimizing the two policies. Timestep $t$ and reward weight terms are omitted for simplicity. The precise expressions for each reward term and their weights are provided in supplementary.
Rewards for Getting Up: $r_{\text{up}}=r_{\text{height}}+r_{\Delta\text{height}}+r_{\text{uprightness}}+r_{\text{stand\_on\_feet}}+r_{\Delta\text{feet\_contact\_forces}}+r_{\text{symmetry}}$, where
• $r_{\text{height}}$ encourages the robot’s height to be close to a target height when standing;
• $r_{\Delta\text{height}}$ encourages the robot to continuously increasing its height;
• $r_{\text{uprightness}}$ encourages the robot to increase the projected gravity on $z$ axis so that the robot will stand upright;
• $r_{\text{stand\_on\_feet}}$ encourages the robot to stand on both feet;
• $r_{\Delta\text{feet\_contact\_forces}}$ encourages the robot to increase contact forces applied to the feet continuously;
• $r_{\text{symmetry}}$ reduces the search space by encouraging (but not requiring) the robot to output bilaterally symmetric actions. Past work [63, 33] employed hard symmetry which improves RL sample efficiency at the cost of limiting robots’ DoFs and generalization. Our soft symmetry reward inherits the benefit but mitigates the limitation.
Rewards for Rolling Over:  $r_{\text{roll}}=r_{\text{gravity}}$, where $r_{\text{gravity}}$ encourages the robot to change its body orientation so that its projected gravity is close to the target projected gravity when lying facing up.
III-B2 Stage II: Deployable Policy
This stage trains policy $\pi$ that will be directly deployed in the real world. Policy $\pi$ is trained to imitate an 8$\times$ slowed-down version of the state trajectories discovered in Stage I, while also respecting strong regularization to ensure Sim2Real transferability. We use the typical control regularization rewards and describe them in supplementary. Below, we describe the tracking reward.
Tracking Rewards: The tracking reward, $r_{\text{tracking}}$, encourages the humanoid to act close to the given motion trajectory derived from the discovered motion. $r_{\text{tracking}}=r_{\text{tracking\_DoF}}+r_{\text{tracking\_body}}$, where
• $r_{\text{tracking\_DoF}}$ encourages the robot to move to the same DoF position as the reference motion, and
• $r_{\text{tracking\_body}}$ encourages the robot to move the body to the same position as the reference. Specifically, $r_{\text{tracking\_body}}$ becomes $r_{\text{head\_height}}$ and $r_{\text{head\_gravity}}$ which encourage the robot to track the height and the projected gravity of the robot’s head for getting-up and rolling-over tasks, respectively.
III-C Stage I to Stage II Curriculum
The design of two-stage policy learning intrinsically constructs a hard-to-easy curriculum [7]. Stage I targets motion discovery in easier settings (weak regularization, no variations in terrain, same starting poses, simpler collision geometry). Once motions have been discovered, Stage II solves the task of making the learned motion deployable and generalizable. As our experiments will show, splitting the work into two phases is crucial for successful learning. Specifically, complexity increases from Stage I to Stage II in the following ways:
III-C1 Collision mesh
As shown in Fig. 2, Stage I uses a simplified collision mesh for faster motion discovery, while Stage II uses the full mesh for improved Sim2Real performance.
III-C2 Posture randomization
Stage I learns to get up (and roll over) from a canonical pose, accelerating learning, while Stage II starts from arbitrary initial poses, enhancing generalization. To further speed up Stage I, we mix in standing poses. For Stage II, we generate a dataset $\mathcal{P}$ of 20K supine poses $\mathcal{P}_{\text{supine}}$ and 20K prone poses $\mathcal{P}_{\text{prone}}$ by randomizing initial DoFs from canonical lying poses, dropping the humanoid from 0.5m, and simulating for 10s to resolve self-collisions. We use 10K poses from each set for training and the rest for evaluation.
III-C3 Control Regularization and Terrain Randomization
For Sim2Real transfer, we use the following control regularization terms and environment randomization in Stage II:
• Weak $\rightarrow$ strong control regularization. Weak control regularization in Stage I enables discovery of getting-up / rolling-over motion, while strong control regularization (via smoothness rewards, DoF velocity penalties, etc, see the full list in supplementary) in Stage II encourages more deployable action.
• Fast $\rightarrow$ slow motion speed. Without strong control regularization, Stage I discovers a fast but unsafe getting-up motion ($<$1s), infeasible for real-world deployment. To address this, we slow it to 8s via interpolation, providing stable tracking targets for Stage II, which better aligns with its control regularization.
• Fixed $\rightarrow$ random dynamics and domain parameters. Stage II also employs domain and dynamics randomization via terrain randomization and noise injection. Such randomization has been shown to play a vital role in successful Sim2Real [68].
TABLE I: Simulation results. We compare HumanUP with several baselines on the held-out split of our curated posture set $\mathcal{P}_{\text{supine}}$ and $\mathcal{P}_{\text{prone}}$ using full URDF. All methods are trained on the training split of our posture set $\mathcal{P}$, except for methods HumanUP w/o Stage II and w/o posture randomization. HumanUP solves task ❸ by solving task ❷ and task ❶ consecutively. We do not include the results of baseline 6 (HumanUP w/o Two-Stage Learning) as it cannot solve the task. Tao et al. [65] is trained to directly solving task ❸ as it does not have a rolling over policy. Sim2Real column indicates whether the method can transfer to the real world successfully. We tested all methods in the real world for which the Smoothness and Safety metrics are reasonable, and Sim2Real is false if deployment wasn’t successful. Metrics are introduced in Section V-C.
Sim2Real Task Smoothness Safety
[…]
Getting Up from Supine Poses
Tao et al. [65] 92.62 $\pm$ 0.54 1.27 $\pm$ 0.00 5.39 $\pm$ 0.01 0.48 $\pm$ 0.00 650.19 $\pm$1.26 0.72 $\pm$ 3.10e-4 0.73 $\pm$ 1.39e-4
HumanUP w/o Stage II 24.82 $\pm$ 0.25 0.83 $\pm$ 0.00 13.70 $\pm$ 0.18 0.71 $\pm$ 0.00 1311.22 $\pm$ 8.57 0.57 $\pm$ 1.45e-3 0.67 $\pm$ 5.56e-4
HumanUP w/o Full URDF 93.95 $\pm$ 0.24 1.22 $\pm$ 0.00 0.71 $\pm$ 0.00 0.11 $\pm$ 0.00 104.14 $\pm$ 0.57 0.92 $\pm$ 8.36e-5 0.77 $\pm$ 9.40e-5
HumanUP w/o Posture Rand. 65.39 $\pm$ 0.50 1.09 $\pm$ 0.04 0.75 $\pm$ 0.05 0.15 $\pm$ 0.03 141.52 $\pm$ 0.61 0.91 $\pm$ 2.32e-4 0.74 $\pm$ 7.24e-5
HumanUP w/ Hard Symmetry 84.56 $\pm$ 0.11 1.23 $\pm$ 0.00 0.97 $\pm$ 0.01 0.22 $\pm$ 0.00 182.39 $\pm$ 0.22 0.89 $\pm$ 1.70e-5 0.78 $\pm$ 8.81e-5
HumanUP 95.34 $\pm$ 0.12 1.24 $\pm$ 0.00 0.56 $\pm$ 0.01 0.10 $\pm$ 0.00 91.74 $\pm$ 0.33 0.93 $\pm$ 1.55e-5 0.78 $\pm$ 4.15e-5
Rolling Over from Prone to Supine Poses
HumanUP w/o Stage II 43.48 $\pm$ 0.41 0.91 $\pm$ 0.00 3.32 $\pm$ 0.31 0.40 $\pm$ 0.05 1684.66 $\pm$ 0.43 0.65 $\pm$ 6.28e-4 0.72 $\pm$ 7.18e-5
HumanUP w/o Full URDF 87.73 $\pm$ 0.33 0.97 $\pm$ 0.00 0.33 $\pm$ 0.00 0.07 $\pm$ 0.00 59.01 $\pm$ 0.05 0.93 $\pm$ 7.91e-5 0.75 $\pm$ 9.98e-5
HumanUP w/o Posture Rand. 37.27 $\pm$ 1.14 0.77 $\pm$ 0.01 0.77 $\pm$ 0.01 0.15 $\pm$ 0.00 234.46 $\pm$ 1.00 0.90$\pm$ 4.98e-4 0.72 $\pm$ 2.04e-4
HumanUP w/ Hard Symmetry 75.53 $\pm$ 0.25 0.60 $\pm$ 0.00 0.31 $\pm$ 0.00 0.09 $\pm$ 0.00 84.95 $\pm$ 0.33 0.95 $\pm$ 3.12e-5 0.76 $\pm$ 2.49e-5
HumanUP 94.40 $\pm$ 0.21 0.99 $\pm$0.00 0.31 $\pm$ 0.00 0.06 $\pm$ 0.00 57.08 $\pm$ 0.20 0.95 $\pm$ 1.51e-4 0.76 $\pm$ 2.48e-5
Getting Up from Prone Poses
Tao et al. [65] 98.99 $\pm$ 0.20 1.26 $\pm$ 0.00 11.73 $\pm$ 0.01 0.76 $\pm$ 0.00 1015.27 $\pm$ 0.65 0.67 $\pm$ 2.24e-4 0.68 $\pm$ 6.41e-5
HumanUP w/o Stage II 27.59 $\pm$ 0.28 1.23 $\pm$ 0.00 5.56 $\pm$ 0.36 0.45 $\pm$ 0.04 1213.07 $\pm$ 5.56 0.67 $\pm$ 4.71e-3 0.71 $\pm$ 2.17e-3
HumanUP w/o Full URDF 89.59 $\pm$ 0.29 0.82 $\pm$ 0.00 0.44 $\pm$ 0.01 0.08 $\pm$ 0.00 77.61 $\pm$ 0.86 0.92 $\pm$ 2.88e-5 0.75 $\pm$ 3.19e-5
HumanUP w/o Posture Rand. 30.25 $\pm$ 0.24 0.87 $\pm$ 0.02 1.05 $\pm$ 0.01 0.15 $\pm$ 0.00 208.23 $\pm$ 1.27 0.90 $\pm$ 3.06e-4 0.73 $\pm$ 1.01e-4
HumanUP w/ Hard Symmetry 67.12 $\pm$ 0.34 1.09 $\pm$ 0.01 0.94 $\pm$ 0.01 0.23 $\pm$ 0.01 196.17 $\pm$ 3.68 0.91 $\pm$ 3.54e-5 0.76 $\pm$ 4.45e-5
HumanUP 92.10 $\pm$ 0.46 1.24 $\pm$ 0.00 0.39 $\pm$ 0.01 0.07 $\pm$ 0.00 69.98 $\pm$ 0.45 0.94 $\pm$ 1.82e-4 0.77 $\pm$ 3.70e-4
IV Implementation Details
IV-A Platform Configurations
We use the Unitree G1 platform [70] in all real-world and simulation experiments. G1 is a medium-sized humanoid robot with 29 actuatable degrees of freedom (DoF) in total. Specifically, the upper body has 14 DoFs, the lower body has 12 DoFs, and the waist has 3 DoFs. As getting up does not involve object manipulation, we disable the 3 DOFs in the wrists, resulting in 23 DoFs in total. Unlike previous robots, G1 has waist yaw and roll DoFs, and we find them useful for our getting-up task. The robot has an IMU sensor for roll and pitch states, and the joint states can be obtained from the motor encoders. We use position control where the torque is derived by a PD controller operating at 50 Hz.
IV-B Simulation Configurations
We use Isaac Gym [52] for simulated training and evaluation. We use a URDF with simplified collision for stage-I training and the official whole-body URDF from Unitree [70] when training stage-II. To accurately model the numerous contacts between the humanoid and the ground, we use a high simulation frequency of 1000 Hz, while the low-level PD controller frequency operates at 50 Hz.
V Simulation Results
V-A Tasks
We evaluate three tasks involved in the humanoid getting-up process: ❶ getting up from supine poses, ❷ rolling over from prone to supine poses, and ❸ getting up from prone poses which can be addressed by solving task ❷ and task ❶ consecutively. Simulation tests are conducted with the full URDF.
V-B Baselines
We compare to the following baselines,
  • a)
    RL with Simple Task Rewards (Tao et al. [65]): This policy is trained with RL using rewards from Tao et al. [65] originally designed for physically animated characters instead of humanoid robots. Similar to our method, this baseline applies a three-stage strong-to-weak torque limit and motion speed curriculum for getting-up policy learning. Because [65] does not consider sim2real deployment regularization and requirements (e.g., smoothness and collision mesh usage), policies learned through their scheme aren’t appropriate for real-world humanoid deployment.
  • b)
    HumanUP w/o Stage II: Our policy trained with only stage I, where no deployment constraints are applied.
  • c)
    HumanUP w/o Full URDF: Our policy trained with two stages, but stage II uses the simplified collision mesh.
  • d)
    HumanUP w/o Posture Randomization: Our policy trained on a single canonical lying posture without any randomization of initialization postures.
  • e)
    HumanUP w/ Hard Symmetry: Our policy trained using a humanoid with a symmetric controller. This symmetric controller follows the symmetry control principle of the G1 controller baseline described in real-world experiments, which leads to bilaterally symmetric control. We set all pitch DoFs actions to be the same between the left and the right DoFs, while flipping the directions of all the roll and yaw actions.
  • f)
    HumanUP w/o Two-Stage Learning: Our policy trained in a single stage with the full collision mesh, posture randomization, and all rewards and regularization terms applied at the same time.
V-C Metrics
• Task Success. We report task success rate Success (%) and Task Metrics: the head height (m) for the getting-up task, and cosine of the angle between the robot’s X-axis (sticking out to the front from the torso) and the gravity vector, for the rolling over task.
• Smoothness. We measure the Action Jitter (rad/$\text{s}^{3}$), DoF Pos Jitter (rad/$\text{s}^{3}$), and mean Energy ($N\cdot$ rad/s) for action smoothness evaluation. The jitter metrics are computed as the third derivative values [16], which indicates the coordination stability of body movements.
• Safety. We introduce safety scores $\mathcal{S}^{\text{Torque}}_{\delta,\alpha}\in[0,1]$ and $\mathcal{S}^{\text{DoF}}_{\delta,\beta}\in[0,1]$ that measure the relative magnitude of commanded torque / DoF compared to the torque and DoF limits, where $\delta$ is a safety threshold. This is essential for robotic safety during execution, as large torques or DoFs will lead to overheating issues and cause mechanical and motor damage. Formally, these scores are defined as:
$\mathcal{S}^{\text{Torque}}_{\delta,\alpha}$ $=1-\Bigg{(}\frac{\alpha}{TJ}\sum_{t,j}\frac{|\tau_{t,j}|}{\tau_{j}^{\max}}+\frac{1-\alpha}{TJ}\sum_{t,j}\mathbbm{1}\Big{(}\frac{|\tau_{t,j}|}{\tau_{j}^{\max}}>\delta\Big{)}\Bigg{)},$
$\mathcal{S}^{\text{DoF}}_{\delta,\beta}=1-\Bigg{(}\frac{\beta}{TJ}\sum_{t,j}\frac{|q_{t,j}|}{q_{j}^{\max}}+\frac{1-\beta}{TJ}\sum_{t,j}\mathbbm{1}\Big{(}\frac{|q_{t,j}|}{q_{j}^{\max}}>\delta\Big{)}\Bigg{)},$
where $\tau_{t,j}$ and $q_{t,j}$ denote the applied torque and joint displacement at time step $t$ for joint $j$, respectively. $\tau_{j}^{\max}$ and $q_{j}^{\max}$ represent their respective limits, $T$ is the total number of time steps, and $J$ is the total number of joints. The threshold $\delta$ determines when a torque or displacement is considered excessive. The indicator function $\mathbbm{1}(\cdot)$ returns 1 if the condition is met and 0 otherwise. The parameters $\alpha,\beta\in[0,1]$ control the trade-off between peak and prolonged violations, ensuring a balanced assessment of safety risks. In this paper, we use $\delta=0.8$, $\alpha=0.5$, $\beta=0.5$ as default during evaluation.
V-D Results and Analysis
Table I presents evaluation results based on policies tested on held-out subsets of our curated initial posture set $\mathcal{P}$—10K samples each from $\mathcal{P}_{\text{supine}}$ and $\mathcal{P}_{\text{prone}}$. Fig. 4 shows the learning curve for the getting-up task, where the termination base height reflects the robot’s ability to lift its body, and body uprightness indicates whether it achieves a stable standing posture.
Figure 3: Real-world results. We evaluate HumanUP (ours) in several real-world setups that span diverse surface properties, including both man-made and natural surfaces, and cover a wide range of roughness (rough concrete to slippery snow), bumpiness (flat concrete to tiles), ground compliance (completely firm concrete to being swampy muddy grass), and slope (flat to about $10^{\circ}$). We compare HumanUP with G1’s built-in getting-up controller and our HumanUP w/o posture randomization (PR). HumanUP succeeds more consistently (78.3% vs 41.7%) and can solve terrains that the G1’s controller can’t.
V-D1 Ignoring Torque / Control Limits Leads to Undeployable Policies
While [65] and HumanUP achieve similar success rates, the smoothness and safety metrics for [65] are significantly worse than HumanUP. For example, the average action jitter metric is nearly 18$\times$ higher than HumanUP. Actions from [65] are highly unstable and unsafe and thus cannot be safely deployed to the real robot. Furthermore, [65] learns a very fast getting-up motion that keeps jumping after getting up. See visualization [65]’s getting up motion in the supplementary material. A similar trend can be seen when comparing HumanUP to HumanUP w/o Stage II. While HumanUP w/o Stage II also solves the task to some extent, it achieves unsatisfying smoothness and safety metrics similar making it inappropriate for real-world deployment. Thus, the regularization imposed in Stage II is essential to making policies more amenable to Sim2Real transfer.
V-D2 Importance of Learning via a Curriculum
So, while it is clear that we need to incorporate strong control regularization for good safety metrics and sim2real transfer, our 2 stage process is better than doing it in a single stage. In fact, as plotted in Fig. 4, HumanUP w/o Two Stage Learning where the policy is trained in a single stage using all sim2real regularization fails to solve the task. This is because the strict Sim2Real regularization makes task learning extremely challenging. Our two-stage curriculum successfully incorporates both aspects: it learns to solve the task, but the policy also operates safely.
V-D3 Full URDF vs. Simplified URDF
Somewhat surprisingly, even though HumanUP w/o Full URDF was trained without the full URDF mesh, it generalizes fine when tested with the full URDF in simulation, as reported in Table I. However, we found poor transfer of this policy to the real world. It failed on all 5 trials on a simple flat terrain. We believe the poor real-world performance was because of the mismatch between the contact it was expecting and the contact that actually happened.
V-D4 Posture randomization helps
HumanUP w/o posture randomization (PR) works much worse than HumanUP suggesting that PR is necessary for generalizable control.
V-D5 Soft symmetry vs. hard symmetry
Compared to HumanUP w/ Hard Symmetry, HumanUP achieves better task success in Table I, particularly for the rolling-over task which is very difficult with symmetric commands.
Figure 4: Learning curve. (a) Termination height of the torso, indicating whether the robot can lift the body. (b) Body uprightness, computed as the projected gravity on the $z$-axis, normalized to $[0,1]$ for better comparison. The overall number of simulation sampling steps is about 5B, normalized to $[0,1]$.
Figure 5: Getting up comparison with G1 controller. G1 controller uses a handcrafted motion trajectory, which can be divided into three phases, while our HumanUP learns a continuous and more efficient whole-body getting-up motion. Our HumanUP enables the humanoid to get up within 6 seconds, half of the G1 controller’s 11 seconds of control. (a), (b), and (c) record the corresponding mean motor temperature of the upper body, lower body, and waist, respectively. G1’s default controller’s execution causes the arm motors to heat up significantly, whereas our policy makes more use of the leg motors that are larger (higher torque limit of 83N as opposed to 25N for the arm motors) and thus able to take more load.
Figure 6: Qualitative examples of failure modes on grass slope and snow field. G1 controller isn’t able to squat on the sloping grass and slips on the slow. HumanUP policy is able to partially get up on both the slope and the snow but falls due to unstable feet placement on the slope and slippage on the snow.
VI Real World Results
We also tested HumanUP policies in the real world on G1 robot. Our real-world test bed consists of 6 different terrains shown in Fig. 3: concrete, brick, stone tiles, muddy grass, grassy slope, and a snow field. These terrains span diverse surface properties, including both man-made and natural surfaces, and cover a wide range of roughness (rough concrete to slippery snow), bumpiness (flat concrete to tiles), ground compliance (completely firm concrete to being swampy muddy grasp), and slope (flat to $\sim 10^{\circ}$). We tested two tasks: a) getting up from supine poses, and b) rolling over from prone to supine poses.
We compare our final HumanUP policy with 1) G1 controller and 2) a high-performing ablation of HumanUP (HumanUP w/o posture randomization). The G1 controller, that comes with the robot, tracks a hand-crafted trajectory in three-phase shown in Fig. 5: Phase 0 brings the robot to a canonical lying pose; Phase 1 first props up and then slides the torso forward using hands, followed by bending legs to squat; Phase 2 uses its waist to tilt up the torso to stand up from squatting. The motions in phase 1 and phase 2 are symmetric, and this default G1 controller only works for supine poses.
VI-A Results
Fig. 3 presents the experimental results. Overall we find that HumanUP policies perform better than the G1 controller and the version of HumanUP without posture randomization (PR). We discuss the results and observed behavior further.
VI-A1 Getting up from supine poses
The default G1 controller works under nominal conditions, i.e., firm, flat concrete ground with a reasonable friction value. However, it starts to fail on more challenging terrains. For the bumpier and rougher terrains (brick surface and stone tiles), the arms may get stuck between bumps causing failures. On slopes, the robot fails to squat or hoist itself up due to both the resistance of the grassland and the unstable squatting pose prone to falling caused by slopes. On the compliant ground, the robot gets destabilized due to the compliance of the ground. On slippery snow, the robot slips.
Both versions of HumanUP outperform the default G1 controller. Trained with terrain and domain randomization, they are robust to real-world variations such as slipperiness, bumps, and slopes. Dynamics randomization further enhances resilience to minor perturbations like slippage or ground compliance. The full method, incorporating posture randomization, performs better than the variant without it, as it is specifically trained to handle diverse initial configurations. Overall, HumanUP achieves a 78.3% success rate in getting up.
VI-A2 Rolling over from prone to supine poses
Findings are similar for the rolling over task. As noted, the default controller can’t handle this situation. The full model exhibits more robust performance than the model trained without posture randomization. Rolling over seems to be easier than getting up, HumanUP achieves a 98.3% success rate for rolling over.
VI-B Motion analysis
Fig. 5 show the motion and recorded motor temperatures for executions of the G1 controller and HumanUP policy.
VI-B1 Motor temperature
Even when the G1 controller works, it uses the arms during Phase 1 of getting up. Fig. 5 (a) shows that the default controller’s execution causes the arm motors to heat up significantly more when compared to HumanUP execution. In contrast, our policy makes more use of the leg motors that are larger (higher torque limit of $83N$ as opposed to $25N$ for the arm motors) and thus able to take more load.
VI-B2 Efficiency
In addition, we find that HumanUP gets the robot to stand successfully within about 6 seconds through a smooth and continuous motion, which is over 2$\times$ more efficient compared to the G1 controller which takes nearly 11 seconds.
VI-C Failure Mode Analysis
Fig. 6 shows example failure modes for the G1 controller and HumanUP on challenging terrains. Fig. 6 (a) shows that the G1 controller tries to utilize the robot’s hands to squat, while the sloping ground prevents it from getting to the full squatting pose due to high friction and weak waist torques to move against the dumping tendency. In contrast, HumanUP manages to lift the body, while the sloping ground sometimes causes an unstable foot orientation. Fig. 6 (b) shows that on even more challenging terrains like snow fields, both G1 and HumanUP controllers may fail due to the slippery and deformable ground.
VII Limitations
HumanUP has several limitations: 1) It relies on high-performance physics platforms like IsaacGym [52] to simulate contact-rich tasks such as getting up and rolling over. However, current robotics simulators lag behind those in animation and gaming, limiting the accuracy and efficiency of contact dynamics simulation, advancements like Genesis [6] and Mujoco Playground [75] are helpful. 2) The RL process in HumanUP is a under-specified problem [3, 13], making it difficult to ensure that the learned motion aligns precisely with human-like behavior. e.g., our discovered motion tends to raise hands for balance.
VIII Discussion
In this paper, we tackle the problem of learning getting-up controllers for real-world humanoid robots. Different from locomotion tasks, getting up involves complex contact patterns that are not known apriori. We develop a two-stage solution for this problem based on reinforcement learning and sim-to-real. Stage I finds a solution under minimal constraints, while Stage II learns to track the trajectory discovered in Stage I under regularization on control and from varied starting poses and on varied terrains. We found this two stage strategy to be effective both in simulation and the real world. Specifically, it enabled us to get a real-world G1 humanoid to stand up from a supine pose and roll over from a supine pose to a prone pose on different terrains and from different starting poses. HumanUP achieves a higher success rate than the default G1 controller and expands the capabilities of the G1 robot.
We hope our learned policies for automatic fall recovery will be useful to researchers and practitioners, while our two stage learning framework may be helpful for other problems that require figuring out complex contact patterns.
Appendix A Material Statements
In our supplementary material, we provide a video demo presenting HumanUP, and this PDF for supplementary information including our reward designs, training details, and additional qualitative analysis.
Rewards
A-1 Rewards components in Stage I
Detailed reward components used in Stage I are summarized in Table II.
TABLE II: Reward components and weights in Stage I. Penalty rewards prevent undesired behaviors for sim-to-real transfer, regularization refines motion, and task rewards ensure successful getting up or rolling over.
Term Expression Weight
Penalty:
Torque limits $\mathds{1}({{\boldsymbol{{\tau}}_{t}}\notin[\boldsymbol{\tau}_{\min},\boldsymbol{\tau}_{\max}]})$0.1
DoF position limits $\mathds{1}({{\boldsymbol{{d}}_{t}}\notin[\boldsymbol{q}_{\min},\boldsymbol{q}_{\max}]})$5
Energy $\lVert\ \boldsymbol{\tau}\odot\dot{\mathbf{q}}\rVert$1e-4
Termination $\mathds{1}_{\text{termination}}$500
Regularization:
DoF acceleration $\lVert{\boldsymbol{\ddot{d}}_{t}}\rVert_{2}$1e-7
DoF velocity $\lVert{\boldsymbol{\dot{d}}_{t}}\rVert_{2}^{2}$1e-4
Action rate $\lVert\boldsymbol{a}_{t}\rVert_{2}^{2}$0.1
Torque $\lVert{\boldsymbol{{\tau}}_{t}}\rVert$6e-7
DoF position error $\mathds{1}(h_{\text{base}}\geq 0.8)\cdot\exp\left(-0.05\|{\boldsymbol{{d}}_{t}}-{\boldsymbol{{d}}_{t}}^{\text{default}}|\right)$0.75
Angular velocity $\lVert\omega^{2}\rVert$0.1
Base velocity $\lVert\boldsymbol{v}^{2}\rVert$0.1
Foot slip $\mathds{1}(\boldsymbol{F}^{\text{feet}}_{z}>5.0)\cdot\sqrt{\|\boldsymbol{v}_{z}^{\text{feet}}\|}$1
Getting-Up Task Rewards:
Base height exp $\exp(\boldsymbol{h}^{base})-1$ 5
Head height exp $\exp(\boldsymbol{h}^{head})-1$ 5
$\Delta$ base height $\mathds{1}(\boldsymbol{h}_{t}^{base}>\boldsymbol{h}^{base}_{t-1})$ 1
Feet contact forces reward $\mathds{1}(\lVert\boldsymbol{F}_{t}^{feet}\rVert>\lVert\boldsymbol{F}^{\text{feet}}_{t-1})\rVert$ 1
Standing on feet reward $\mathds{1}\big{(}(\lVert\boldsymbol{F}^{\text{feet}}\rVert>0\big{)}\&\big{(}\boldsymbol{h}^{\text{feet}}<0.2)\big{)}$ 2.5
Body upright reward $\exp(-\mathbf{g}^{\text{base}}_{z})$ 0.25
Feet height reward $\exp(-10\cdot\boldsymbol{h}^{\text{feet}})$ 2.5
Feet distance reward $\frac{1}{2}\Big{(}\exp(-100\left|\max(\boldsymbol{d}_{\text{feet}}-\boldsymbol{d}_{\min},-0.5)\right|)$ $+\exp(-100\left|\max(\boldsymbol{d}_{\text{feet}}-\boldsymbol{d}_{\max},0)\right|)\Big{)}$ 2
Foot orientation $\sqrt{\lVert G_{xy}^{\text{feet}}\rVert}$0.5
Soft body symmetry penalty $\left\|\mathbf{a}_{\text{left}}-\mathbf{a}_{\text{right}}\right\|$1.0
Soft waist symmetry penalty $\left\|\mathbf{a}^{\text{waist}}\right\|$1.0
Rolling-Over Task Rewards:
Base Gravity Exp $\frac{1}{2}\Big{(}\exp\big{(}-0.01(1-\cos\theta_{\text{left}})\big{)}+$ $\exp\big{(}-0.01(1-\cos\theta_{\text{right}})\big{)}\Big{)},$ $\cos\theta=\frac{\mathbf{g}^{\text{knee}}\cdot\mathbf{g}_{\text{target}}}{\|\mathbf{g}^{\text{base}}\|\|\mathbf{g}_{\text{base}}\|}$ 8
Knee Gravity Exp $\frac{1}{2}\Big{(}\exp\big{(}-0.01(1-\cos\theta_{\text{left}})\big{)}+$ $\exp\big{(}-0.01(1-\cos\theta_{\text{right}})\big{)}\Big{)},$ $\cos\theta=\frac{\mathbf{g}^{\text{knee}}\cdot\mathbf{g}_{\text{target}}}{\|\mathbf{g}^{\text{base}}\|\|\mathbf{g}_{\text{base}}\|}$ 8
Feet distance reward $\frac{1}{2}\Big{(}\exp(-100\left|\max(\boldsymbol{d}_{\text{feet}}-\boldsymbol{d}_{\min},-0.5)\right|)$ $+\exp(-100\left|\max(\boldsymbol{d}_{\text{feet}}-\boldsymbol{d}_{\max},0)\right|)\Big{)}$ 2
Feet height reward $\exp(-10\cdot\boldsymbol{h}^{\text{feet}})$ 2.5
A-2 Rewards components in Stage II
Detailed reward components used in Stage II are summarized in Table III.
TABLE III: Reward components and weights in Stage II. Penalty rewards prevent undesired behaviors for sim-to-real transfer, regularization refines motion, and task rewards ensure successful whole-body tracking in real time.
Term Expression Weight
Penalty:
Torque limits $\mathds{1}({{\boldsymbol{{\tau}}_{t}}\notin[\boldsymbol{\tau}_{\min},\boldsymbol{\tau}_{\max}]})$5
Ankle torque limits $\mathds{1}({{\boldsymbol{{\tau}}_{t}}^{\text{ankle}}\notin[\boldsymbol{\tau}^{\text{ankle}}_{\min},\boldsymbol{\tau}^{\text{ankle}}_{\max}]})$0.01
Upper torque limits $\mathds{1}({{\boldsymbol{{\tau}}_{t}}^{\text{upper}}\notin[\boldsymbol{\tau}^{\text{upper}}_{\min},\boldsymbol{\tau}^{\text{upper}}_{\max}]})$0.01
DoF position limits $\mathds{1}({{\boldsymbol{{d}}_{t}}\notin[\boldsymbol{q}_{\min},\boldsymbol{q}_{\max}]})$5
Ankle DoF position limits $\mathds{1}({{\boldsymbol{{d}}_{t}}^{\text{ankle}}\notin[\boldsymbol{q}^{\text{ankle}}_{\min},\boldsymbol{q}^{\text{ankle}}_{\max}]})$5
Upper DoF position limits $\mathds{1}({{\boldsymbol{{d}}_{t}}^{\text{upper}}\notin[\boldsymbol{q}^{\text{upper}}_{\min},\boldsymbol{q}^{\text{upper}}_{\max}]})$5
Energy $\lVert\ \boldsymbol{\tau}\odot\dot{\mathbf{q}}\rVert$1e-4
Termination $\mathds{1}_{\text{termination}}$50
Regularization:
DoF acceleration $\lVert{\boldsymbol{\ddot{d}}_{t}}\rVert_{2}$1e-7
DoF velocity $\lVert{\boldsymbol{\dot{d}}_{t}}\rVert_{2}^{2}$1e-3
Action rate $\lVert\boldsymbol{a}_{t}\rVert_{2}^{2}$0.1
Torque $\lVert{\boldsymbol{{\tau}}_{t}}\rVert$0.003
Ankle torque $\lVert{\boldsymbol{{\tau}}_{t}}^{\text{ankle}}\rVert$6e-7
Upper torque $\lVert{\boldsymbol{{\tau}}_{t}}^{\text{upper}}\rVert$6e-7
Angular velocity $\lVert\omega^{2}\rVert$0.1
Base velocity $\lVert\boldsymbol{v}^{2}\rVert$0.1
Tracking Rewards:
Tracking DoF position $\exp\left(-\frac{({\boldsymbol{{d}}_{t}}-{\boldsymbol{{d}}_{t}}^{\text{target}})^{2}}{4}\right)$ 8
Feet distance reward $\frac{1}{2}\Big{(}\exp(-100\left|\max(\boldsymbol{d}_{\text{feet}}-\boldsymbol{d}_{\min},-0.5)\right|)$ $+\exp(-100\left|\max(\boldsymbol{d}_{feet}-\boldsymbol{d}_{\max},0)\right|)\Big{)}$ 2
Foot orientation $\sqrt{\lVert\mathbf{g}_{xy}^{\text{feet}}\rVert}$0.5
Training Details
HumanUP contains two-stage learning, which are both trained in simulation. In Stage I, we train the discovery policy $f$ for overall 5B simulation steps in total, and 20K simulation steps for Stage II deployable policy $\pi$. All training is conducted on IsaacGym [52], and we train our policies using 4,096 paralleled environments on a single NVIDIA RTX 4090 or L40S GPU. For the getting-up task, we slow down the discovered trajectory to 8 seconds. For the rolling-over task, the trajectory is slowed down to 4 seconds. We use flat terrains in Stage I and varied terrains during Stage II training, involving flat, rough, and slopes. We follow previous works [29, 10, 77] to apply varied dynamics randomization such as friction, base CoM offset, and control delay.
Additional Qualitative Simulation Baseline Results
Fig. 7 showcases a visualization of getting up from a prone pose generated by baseline method [65]. It can be observed that this method generates motion that is highly unstable and unsafe to deploy in the real world. For example, its joints continuously jitter, the feet are stumbling and the body is keep jumping up. This indicates that this baseline [65] cannot be Sim2Real.